nevaehmar:

\(_8C_3\cdot _5C_2\) a. 6720 b. 5.6 c. 3360 d. 560

dude:

For combinatorics, you probably want to use a calculator or you can use the equation \(_nC_r = \large \frac{n!}{( r! (n - r)! )}\) (In case you don't know what ! mean) ! = factorial, multiply all the numbers from that value all the way to 1 5! = 5\(\times\)4\(\times\)3\(\times\)2\(\times\)1